OpenStudy (anonymous):

Through : (2,4), slope = 9/2 Write the slope intercept form of the equation of the line through the given point with the given slope

OpenStudy (texaschic101):

y = mx + b slope(m) = 9/2 (2,4)...x = 2 and y = 4 now sub 4 = 9/2(2) + b 4 = 9 + b 4 - 9 = b -5 = b so your equation is : y = 9/2x - 5

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

Umm how about through (1,-3), slope=0?